# Stability for the 2-D plane Poiseuille flow in finite channel

### 31 Dec 2023 v1

@@ 5. Nonlinear stability
$\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}$ $\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$
$\lesssim$ $\|\overline{\omega}\|_{L^{\infty}L^{2}}\|\omega_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{6}}|k|^{-\frac{1}{3}}\mathcal{E}_{0}\mathcal{E}_{k}.$
@@
$\|(f_{1})_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{6}}|k|^{-\frac{1}{3}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}$ (5.10)
@@
$\mathcal{E}_{k}\lesssim$ $\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{3}}|k|^{\frac{1}{3}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}}$
$\lesssim$ $\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l},$
@@
$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\lesssim\sum_{k\in\mathbb{Z}}\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\sum_{k\in\mathbb{Z}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}.$
Due to $\|u_{0}\|_{H^{\frac{3}{2}+}}\leq\epsilon_{1}\nu^{1/2}$, then one has
$\sum_{k\in\mathbb{Z}}\|\omega_{0,k}\|_{L^{2}}\leq\epsilon_{1}C\nu^{\frac{1}{2}}.$
@@
$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\leq\epsilon_{1}C\nu^{\frac{1}{2}}.$
@@ Acknowledgment.
The authors thank to Prof. Zhifei Zhang for valuable discussions and suggestions. S. Ding is supported by the Key Project of National Natural Science Foundation of China under Grant 12131010, National Natural Science Foundation of China under Grant 12271032 and Natural Science Foundation of Guangdong Province under Grant 2021A1515010249 and 2021A1515010303. Z. Lin is partially supported by National Natural Science Foundation of China under Grant 11971009.

### 03 Mar 2024 v2

@@ 5. Nonlinear stability
[…]
$\lesssim$ $\|\overline{\omega}\|_{L^{\infty}L^{2}}(\nu|k|)^{-\frac{1}{4}\cdot\frac{5}{6}}(\nu^{\frac{1}{2}}|k|)^{-\frac{1}{6}}\mathcal{E}_{k}$
$+\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{k}\|_{L^{2}L^{2}}^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}(\omega)_{k}\|_{L^{2}L^{2}}^{\frac{1}{2}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$
$\lesssim$ $\nu^{-\frac{7}{24}}|k|^{-\frac{3}{8}}\mathcal{E}_{0}\mathcal{E}_{k}.$
@@
$\|e^{c\nu^{\frac{1}{2}}t}(f_{1})_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{7}{24}}|k|^{-\frac{3}{8}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}$ (5.5)
@@
$\mathcal{E}_{k}\lesssim$ $\|\Delta_{k}\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{8}}|k|^{\frac{3}{8}}\|e^{c\nu^{\frac{1}{2}}t}(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}(f_{2})_{k}\|_{L^{2}L^{2}}$
$\lesssim$ $\|\Delta_{k}\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{2}{3}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l},$
@@
$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\lesssim\sum_{k\in\mathbb{Z}}\|\Delta_{k}\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{2}{3}}\sum_{k\in\mathbb{Z}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}.$
Due to $\|u_{0}\|_{H^{\frac{7}{2}+}}\leq\epsilon_{1}\nu^{2/3}$ with
$\sum_{k\in\mathbb{Z}}\|\Delta_{k}\omega_{0,k}\|_{L^{2}}\leq\epsilon_{1}C\nu^{\frac{2}{3}}.$
@@
$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\leq\epsilon_{1}C\nu^{\frac{2}{3}}.$
@@ Acknowledgment.
The authors thank to Prof. Zhifei Zhang for valuable discussions and suggestions. S. Ding is supported by the Key Project of National Natural Science Foundation of China under Grant 12131010, National Natural Science Foundation of China under Grant 12271032. Z. Lin is partially supported by National Natural Science Foundation of China under Grant 11971009.